Hi,
Actually I'm in college algebra, but since the following problem has to do with range, this was the closest category I could come up with.
Here's the problem:
Find the range of the function: f(x)=x^2+12x+34

Found 2 solutions by stanbon, Alan3354:
Answer by stanbon(75887)

Find the range of the function: f(x)=x^2+12x+34
---
Since the parabola opens upward from the vertex,
you have to find the y-value of the vertex.
-----
Vertex occurs at x = -b/2a = -12/(2) = 6
----
y-value at x = 6 is 6^2+12*6+34 = 142
-----
Range: All Real Numbers >= 142
===================================
Cheers,
Stan H.

Answer by Alan3354(69443)

The minimum of the function is the vertex at (-6,-2)
The range is all real numbers >= -2
PS The other tutor lost a minus sign.